A LABORATORY SEARCH FOR THE CARRIER OF U-LINES ATTRIBUTED TO l-C_{3}H^{+} IN THE HORSE-HEAD NEBULA PDR

\underline{MICHAEL C. McCARTHY}, KYLE N. CRABTREE, and OSCAR MARTINEZ, JR., Harvard-Smithsonian Center for Astrophysics, Cambridge, MA 02138.

Recent radio observations of the Horsehead nebula PDR in the millimeter-wave band by Pety et al.^{a} detected a series of unidentified lines which have been attributed to l-C_{3}H^{+}, a molecular cation for which no high-resolution laboratory data presently exist. We have detected a pair of rotational lines in the centimeter-wave band at frequencies predicted from their derived spectroscopic constants using Fourier transform microwave and microwave-microwave double resonance spectroscopies. Experimental evidence strongly suggests that the laboratory and astronomical lines arise from a common carrier, and that the carrier is a closed-shell, reactive hydrocarbon containing a linear or nearly-linear three carbon atom backbone.

aPety et al. Astron. & Astrophys. 548, A68 (2012).